What is the resistivity with the absolute error?

I'm using the formula resistivity = R*A/L

R = 126.32
Answer #1

del (resistivity)/resistivity = del(R)/R + del(A)/A +del(L)/L


resistivity = RA/L = 3.46*10^-3

del(resistivty)/resistivity = 1.97/ 123.32 + 0.05/2.5 + 0.028/9.123 = 0.039

error = del(resistivity) =0.135*10^-3